The following is the number of minutes to commute from home to work for a group of 25 automobile executives 30 32 22 3118 22 42 2830 19 24 38 25 28 37 32 36 36 28 33 27 25 35 29 31 a. How many classes would you recommend? Number of classes This is a numeric cell, so please enter numbers only. b. What class interval would you suggest? (Round up your answer to the next whole number.) Class interval c. Organize the data and plot a frequency distribution on a piece of paper. Comment on the shape of the frequency distribution. It is not symmetric. It is fairly symmetric, with most of the values between 18 and 42.
Answer #1

a).

As we know that the desire numbers of classes are “5” to “20”. So, here the minimum and the maximum numbers are given by “42” and “18”, => the “range” is “42-18=24”, => “24/5 = 4.8. Now, if the number of classes is “6” then “24/6 = 4 (integer)”. SO, here the numbers of classes should be “6”.

b).

So, here the “class intervals” are given by, “(15-19), “(20-24)”, “(25-29)”, “(30-34)”, “(35-39)”, “(40-45)”.

Consider the following table.

Class Interval Class Boundaries Numbers less than 15less than 15 15 - 19 20 24 25 29 30 34 35 39 40- 45 More than 45 Mor

c).

Consider the following frequency distribution.

Class Interval less than 15 15-19 20 24 25- 29 30 34 35- 39 40-45 More than 45 Frequency 0 6 4 0 Series1 less than 15-19 2

So, here we can see that the data is not symmetric in nature.